OpenStudy (anonymous):

I was asked to find the standard form of the equation of this hyperbola 9x^2-y^2 + 54x + 10y + 55 = 0 but as result I got 9(x+3)^2 - (y+5)^2 = 1 and I don't know what to do with the 9...

OpenStudy (amistre64):

you completed your squares?

OpenStudy (amistre64):

9x^2 -y^2 + 54x + 10y + 55 = 0 9x^2 + 54x -y^2 + 10y + 55 = 0 9(x^2 + 6x) -(y^2 -10y) + 55 = 0 9(x^2 + 6x+ 9-9) -(y^2 -10y +25-25) + 55 = 0 9((x+3)^2-9) -((y-5)^2-25) + 55 = 0 9(x+3)^2 -9(9) -(y-5)^2 -(-25) + 55 = 0 9(x+3)^2 -81 -(y-5)^2 +25 + 55 = 0 9(x+3)^2 -81 -(y-5)^2 +80 = 0 9(x+3)^2 -(y-5)^2 = 1

OpenStudy (amistre64):

yeah, thats all there is to it

OpenStudy (anonymous):

Thanks, I think that in order to have the equation in standard form I need to (x+3)^2 - (y+5)^2/9 = 1/9 (x+3)^2/1/2 - (y+5)^2 = 1

OpenStudy (anonymous):

Sorry (x+3)^2/1/9 *

OpenStudy (amistre64):

no, you have it fine as is

OpenStudy (amistre64):

its equal to 1, you cant divide off the 9 and still have it =1 and having (x-__)^2 ----- (1/9) just isnt proper that im aware of

OpenStudy (amistre64):

if we had something like =6, we would have divided off the 6 to get it into standard form 9/6 = 3/2 and 1/6 is well 1/6 other than that